# Exploring Consistency in Graph Representations: from Graph Kernels to Graph Neural Networks

### v1

@@ 1 Introduction
Graph kernel methods, on the other hand, are designed to capture similarities between graphs and utilize these similarities for classification tasks. For instance, subgraph-pattern approaches (Shervashidze et al., 2009; Costa and Grave, 2010; Kriege et al., 2020) compare graphs by counting the occurrences of fixed-size subgraph motifs. Other methods compare sequences of vertices or edges encountered during graph traversals (Borgwardt and Kriegel, 2005; Kashima et al., 2003; Zhang et al., 2018). Among all graph kernels, two notable ones are the Weisfeiler-Lehman subtree (WL-subtree) kernel (Shervashidze et al., 2011) and the Weisfeiler-Lehman optimal assignment (WLOA) kernel (Kriege et al., 2016). They are found to have comparable performance to simple GNNs (Nikolentzos et al., 2021). The WL-subtree kernel iteratively relabels graphs using the Weisfeiler-Lehman algorithm (Weisfeiler and Lehman, 1968) and constructs a kernel based on the number of occurrences of each label. The WLOA kernel uses the same relabeling scheme but computes a matching between substructures to reveal structural correspondences between the graphs.
@@
In this work, we aim to bridge the gap between kernel methods and GNN models. Given the iterative nature of GNNs, we study a class of kernels which are induced from graph representations obtained through an iterative process and name them iterative graph kernels (IGK). Within this framework, we define the consistency property, which ensures that similar graphs remain similar in subsequent iterations. Our analysis demonstrates that kernels with this property yield better classification performance. Furthermore, we find that this property sheds light on why the WLOA kernel outperforms the WL-subtree kernel. The WLOA kernel asymptotically demonstrates consistency as the iteration goes to infinity, whereas the WL-subtree kernel does not exhibit this behavior. Inspired by these findings and the analogy between message-passing GNNs and the WL-subtree kernel (Xu et al., 2019), we hypothesize that this principle is also applicable to GNNs. To explore this, we introduce a novel loss function designed to align the ranking of graph similarities across GNN layers. The aim is to ensure that the relational structures of the graphs are preserved and consistently reflected throughout the representation space of these layers. We validate this hypothesis by applying our loss function to different GNN backbones across various graph datasets. Extensive experiments demonstrate that our proposed model-agnostic consistency loss improves graph classification performance comprehensively.
@@ 3.2 Key Properties: Monotonic Decrease & Order Consistency
To effectively capture relational structures, we design the IGKs to progressively differentiate between graphs. With each iteration, additional structural features are considered, enabling the distinction of graphs that may have been indistinguishable in earlier iterations. This implies two properties: (1) the kernel values monotonically decrease with larger iterations, as the similarity between two graphs decreases with the consideration of more features; and (2) the similarity rankings across different iterations should remain consistent, meaning that graphs deemed dissimilar in early iterations should not be considered similar in later iterations. We then formally define these two properties and demonstrate how they can lead to a non-decreasing margin (better performance) in the binary classification task.
@@
Next we show that these two properties can lead to a non-decreasing margin in the binary classification task, which suggests better performance.
@@
Let $\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)$ be a normalized iterative graph kernel that is monotonically decreasing and preserves order consistency. In the binary graph classification task with uniform graph representations, the margin between two classes in the representation space is non-decreasing w.r.t the iteration $i$, where the margin at iteration $i$ is defined as the shortest distance between two graph representations from different classes: margin=$\lVert\psi^{i}(x)-\psi^{i}(y)\rVert$, $\mathcal{Y}(x)\neq\mathcal{Y}(y)$.
Proof sketch. Suppose that at iteration $i$, the margin is defined by a pair of graphs, $x_{1}$ from class 1 and $y_{1}$ from class 2. In the next iteration, $i+1$, the margin is determined by another pair, $x_{2}$ and $y_{2}$. Two possibilities arise: (1) $x_{1}=x_{2}$ and $y_{1}=y_{2}$, or (2) $x_{1}\neq x_{2}$ or $y_{1}\neq y_{2}$.
@@
    If the pair defining the margin remains unchanged, the margin at iteration $i+1$ can only increase or stay the same, as the kernel is a monotonically decreasing function, indicating a reduction in similarity.
@@
    We prove case 2 by contradiction, assuming the margin decreases at iteration $i+1$. This would imply that the kernel value for $x_{2}$ and $y_{2}$ increases at iteration $i+1$, which contradicts the fact that the kernel function is monotonically decreasing.
Thus, in both cases, the margin does not decrease as the iterations progress.
We provide the detailed proof in Appendix A.1
@@ 3.3 Theoretical Verification with WL-based Kernels
We include the complete proof in Appendix A.2 and A.3.
@@ 4 Proposed Strategy
Figure 2: Computation of loss. At each layer, pairwise distance matrix $\mathbf{D}$ is calculated using the normalized representations of graphs in a batch. After randomly selecting a reference graph $x_{k}$, the reference probability matrix is computed using the distance matrix from previous layer, where entry $(n,m)$ represents the known probability that the graph $x_{k}$ is more similar to the graph $x_{n}$ than to the graph $x_{m}$. For the distance matrix of current layer, we compute the predicted probability that $x_{k}$ is closer to $x_{n}$ than to $x_{m}$ and form the prediction probability matrix. Consistency loss is computed as the cross-entropy between the predicted and reference probability matrices
@@ 5.1 Experiment Setup
For the TU dataset and Reddit-T, consistent with prior work (Xinyi and Chen, 2019; Xu et al., 2019), we utilize an 8:1:1 ratio for training, validation, and testing sets. For the OGB datasets, we use the official splits provided. Training stops at the highest validation performance, and test accuracy is taken from the corresponding epoch in each fold. Final results are reported as the mean accuracy (except ogbg-molhiv) and standard deviation over $10$ folds. For ogbg-molhiv, we follow the official evaluator and use ROC-AUC as the evaluation metric. Detailed information about these datasets can be found in Appendix B.
@@
We use three widely adopted GNN models as baselines: namely, GCN (Kipf and Welling, 2017), GIN (Xu et al., 2019), and GraphSAGE (Hamilton et al., 2017). We also include two recent GNN models, namely GTransformer (Shi et al., 2021) and GMT (Baek et al., 2021). To ensure a fair comparison, we maintain the same number of layers and layer sizes for both the base models and the models with our proposed consistency loss, ensuring the sharing of the same network architecture. Detailed information about hyperparameter tuning is provided in Appendix C.
@@ 5.2 Effectiveness of Consistency Loss
Table 1: Classification performance on the TU and OGB datasets, with and without the consistency loss. Highlighted cells indicate instances where the base GNN with the consistency loss outperforms the base GNN alone. The reported values are average accuracy for TU datasets and ROC-AUC for the ogbg-molhiv dataset, including their standard deviations.
@@
Furthermore, we analyze the complexity and scalability of our method on the TU Dataset, with additional details provided in Appendices D.1 and D.2. Additionally, we demonstrate the method’s potential to enhance performance significantly, even with a marginal increase in computational cost, as illustrated in Appendix E.
@@ 5.3 Effect of the Consistency Loss on Rank Correlation
Table 2: Spearman correlation was computed for graph representations from consecutive layers on the TU datasets, both with and without consistency loss. Values with higher rank correlation are highlighted in grey. The consistency loss can enhance the rank correlation of graph similarities.
@@ 5.4 Study on Task Complexity
We increase the task complexity by expanding the number of classes that the model needs to classify. To assess the effect of increased class complexity on our method’s performance, we sampled subsets from the REDDIT-MULTI 5K dataset (Yanardag and Vishwanathan, 2015) with a progressively greater number of classes, which originally consists of five classes. Specifically, we randomly sampled between 2 and 4 classes to construct new datasets from the original dataset and conducted classification tasks using both GCN and GCN with $\mathcal{L}_{\text{consistency}}$ on these newly constructed datasets. We report the mean test accuracy over five experiments for each subset, as presented in Table.3.
@@
Furthermore, we evaluate the method’s efficiency under various complexity scenarios, as detailed in the Appendix D.3.
@@ 7 Conclusion
Appendix A Theoretical Verification with WL-based Kernels
A.1 Non-decreasing Margin
Suppose that at iteration $i$, graph $x_{1}$ in class 1 and graph $y_{1}$ in class 2 decide the margin and at iteration $i+1$, graph $x_{2}$ in class 1 and graph $y_{2}$ in class 2 decide the margin.
There are two scenarios: (1) $x_{1}=x_{2}$ and $y_{1}=y_{2}$, (2) $x_{1}\neq x_{2}$ or $y_{1}\neq y_{2}$. For case 1, margin at iteration $i$ is given by:
$\begin{split}\lVert\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1}))\rVert&=\sqrt{(\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1})))^{T}(\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1})))}\\ &=\sqrt{2-2\langle\phi(\psi^{i}(x_{1})),\phi(\psi^{i}(y_{1}))\rangle}=\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{1},y_{1},i)}.\end{split}$ (1)
Since $\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)$ is monotonically decreasing, we have:
$\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{1},y_{1},i)}\leq\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{1},y_{1},i+1)}=\lVert\phi(\psi^{i+1}(x_{1}))-\phi(\psi^{i+1}(y_{1}))\rVert.$ (2)
This suggests that the margin does not decrease.
For case 2, we prove by contradiction and first assume that the margin decreases. Since the margin decreases, we have:
$\lVert\phi(\psi^{i+1}(x_{2}))-\phi(\psi^{i+1}(y_{2}))\rVert<\lVert\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1}))\rVert.$ (3)
At iteration $i$, since $x_{1}$ and $y_{1}$ decide the margin, we have:
$\lVert\phi(\psi^{i}(x_{1}))-\phi(\psi^{i}(y_{1}))\rVert\leq\lVert\phi(\psi^{i}(x_{2}))-\phi(\psi^{i}(y_{2}))\rVert.$ (4)
Combining Equation 3 and Equation 4, we have:
$\lVert\phi(\psi^{i+1}(x_{2}))-\phi(\psi^{i+1}(y_{2}))\rVert<\lVert\phi(\psi^{i}(x_{2}))-\phi(\psi^{i}(y_{2}))\rVert.$ (5)
This is equivalent to:
$\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{2},y_{2},i+1)}<\sqrt{2-2\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x_{2},y_{2},i)},$ (6)

### v2

@@ 1 Introduction
Graph kernel methods, on the other hand, are designed to capture similarities between graphs and utilize these similarities for classification tasks. For instance, subgraph-pattern approaches (Shervashidze et al., 2009; Kriege et al., 2020) compare graphs by counting the occurrences of fixed-size subgraph motifs. Other methods compare sequences of vertices or edges encountered during graph traversals (Borgwardt and Kriegel, 2005; Kashima et al., 2003; Zhang et al., 2018). Among all graph kernels, two notable ones are the Weisfeiler-Lehman subtree (WL-subtree) kernel (Shervashidze et al., 2011) and the Weisfeiler-Lehman optimal assignment (WLOA) kernel (Kriege et al., 2016). They are found to have comparable performance to simple GNNs (Nikolentzos et al., 2021). The WL-subtree kernel iteratively relabels graphs using the Weisfeiler-Lehman algorithm (Weisfeiler and Lehman, 1968) and constructs a kernel based on the number of occurrences of each label. The WLOA kernel uses the same relabeling scheme but computes a matching between substructures to reveal structural correspondences between the graphs.
@@
In this work, we aim to bridge the gap between kernel methods and GNN models. Given the iterative nature of GNNs, we study a class of kernels which are induced from graph representations obtained through an iterative process and name them iterative graph kernels (IGK). Within this framework, we define the consistency property, which ensures that similar graphs remain similar in subsequent iterations. Our analysis demonstrates that kernels with this property ensure better generalization ability, leading to improved classification performance.Furthermore, we find that this property sheds light on why the WLOA kernel outperforms the WL-subtree kernel. The WLOA kernel asymptotically demonstrates consistency as the iteration goes to infinity, whereas the WL-subtree kernel does not exhibit this behavior. Inspired by these findings and the analogy between message-passing GNNs and the WL-subtree kernel (Xu et al., 2019), we hypothesize that this principle is also applicable to GNNs. To explore this, we introduce a novel loss function designed to align the ranking of graph similarities across GNN layers. The aim is to ensure that the relational structures of the graphs are preserved and consistently reflected throughout the representation space of these layers. We validate this hypothesis by applying our loss function to different GNN backbones across various graph datasets. Extensive experiments demonstrate that our proposed model-agnostic consistency loss improves graph classification performance comprehensively.
@@ 3.2 Key Properties: Monotonic Decrease & Order Consistency
To effectively capture relational structures, we design the IGKs to progressively differentiate between graphs. With each iteration, additional structural features are considered, enabling the distinction of graphs that may have been indistinguishable in earlier iterations. This implies two properties: (1) the kernel values monotonically decrease with larger iterations, as the similarity between two graphs decreases with the consideration of more features; and (2) the similarity rankings across different iterations should remain consistent, meaning that graphs deemed dissimilar in early iterations should not be considered similar in later iterations. We then provide formal definitions for these two properties and prove how they can improve generalization in the binary classification task.
@@
Next, we show that these two properties can lead to a monotonically decreasing upper bound on the test error in the binary classification task, which suggests improved performance.
@@
Let $\tilde{\mathbb{K}}_{\mathcal{F}_{c},\phi}(x,y,i)$ be a normalized iterative graph kernel. Its generalization ability on the test set is provably strong when it decreases monotonically and preserves order consistency.
Proof sketch. We consider a binary classification task where supervision is provided through labels and similarity orders. We demonstrate that the presence of these two properties enhances the generalizability of the learned iterative kernel. The proof is structured in two steps:
@@
    We first show that iterative kernels can learn the correct ranking given sufficient data during training, using the PAC-learning framework.
@@
    Next, we demonstrate that, once the correct ranking is learned, an iterative kernel learned with the properties of monotonic decrease and order consistency guarantees that the error rate of the graphs in the test set decreases as the number of iterations increases.
Details of the proof can be found in Appendix A.
@@ 3.3 Theoretical Verification with WL-based Kernels
We include the complete proof in Appendix B.1 and B.2.
@@ 4 Proposed Strategy
Figure 2: Computation of Consistency loss. At each layer, pairwise distance matrix $\mathbf{D}$ is calculated using the normalized representations of graphs in a batch. After randomly selecting a reference graph $x_{k}$, the reference probability matrix is computed using the distance matrix from previous layer, where entry $(n,m)$ represents the known probability that the graph $x_{k}$ is more similar to the graph $x_{n}$ than to the graph $x_{m}$. For the distance matrix of current layer, we compute the predicted probability that $x_{k}$ is closer to $x_{n}$ than to $x_{m}$ and form the prediction probability matrix. Consistency loss is computed as the cross-entropy between the predicted and reference probability matrices
@@ 5.1 Experiment Setup
For the TU dataset and Reddit-T, consistent with prior work (Xinyi and Chen, 2019; Xu et al., 2019), we utilize an 8:1:1 ratio for training, validation, and testing sets. For the OGB datasets, we use the official splits provided. Training stops at the highest validation performance, and test accuracy is taken from the corresponding epoch in each fold. Final results are reported as the mean accuracy (except ogbg-molhiv) and standard deviation over $10$ folds. For ogbg-molhiv, we follow the official evaluator and use ROC-AUC as the evaluation metric. Detailed information about these datasets can be found in Appendix C.
@@
We use three widely adopted GNN models as baselines: namely, GCN (Kipf and Welling, 2017), GIN (Xu et al., 2019), and GraphSAGE (Hamilton et al., 2017). We also include two recent GNN models, namely GTransformer (Shi et al., 2021) and GMT (Baek et al., 2021). To ensure a fair comparison, we maintain the same number of layers and layer sizes for both the base models and the models with our proposed consistency loss, ensuring the sharing of the same network architecture. Detailed information about hyperparameter tuning is provided in Appendix D.
@@ 5.2 Effectiveness of Consistency Loss
Table 1: Classification performance on the TU, OGB and Reddit-T datasets, with and without the consistency loss. Highlighted cells indicate instances where the base GNN with the consistency loss outperforms the base GNN alone. The reported values are average accuracy for TU datasets and ROC-AUC for the ogbg-molhiv dataset, including their standard deviations.
@@
Furthermore, we analyze the complexity and scalability of our method on the TU Dataset, with additional details provided in Appendices E.1 and E.2. Additionally, we demonstrate the method’s potential to enhance performance significantly, even with a marginal increase in computational cost, as illustrated in Appendix F.
@@ 5.3 Effect of the Consistency Loss on Rank Correlation
Table 2: Spearman correlation was computed for graph representations from consecutive layers on the TU datasets, both with and without consistency loss. Values with higher rank correlation are highlighted. The consistency loss can enhance the rank correlation of graph similarities.
@@ 5.4 Study on Task Complexity
We increase the task complexity by expanding the number of classes that the model needs to classify. To assess the effect of increased class complexity on our method’s performance, we sampled subsets from the REDDIT-MULTI 5K dataset (Yanardag and Vishwanathan, 2015) with a progressively greater number of classes, which originally consists of five classes. Specifically, we randomly sampled between 2 and 4 classes to construct new datasets from the original dataset and conducted classification tasks using both GCN and GCN with $\mathcal{L}_{\text{consistency}}$ on these newly constructed datasets. We report the mean test accuracy over five experiments for each subset, as presented in Table 3.
@@
Furthermore, we evaluate the method’s efficiency under various complexity scenarios, as detailed in the Appendix E.3.
@@ 7 Conclusion
Appendix A Proof of Theorem 3.4
A.1 Learning Correct Similarity Order with Iterative Kernels
In this section, we first prove that the iterative kernel with bounded feature mapping is capable of learning the pairwise similarity order given sufficient training data, as shown by the PAC-learning framework. We consider a dataset consisting of instances $x\in\chi$, where $\chi$ is the input space. The dataset is denoted as $\chi=\left\{x_{1},x_{2},\ldots,x_{n}\right\}$. We employ a kernel function $\mathbb{K}:\chi\times\chi\rightarrow\mathbb{R}$ with an associated feature mapping $\phi:\chi\rightarrow\mathcal{H}_{\mathbb{K}}$ such that $\mathbb{K}(x,y)=\langle\phi(x),\phi(y)\rangle_{\mathcal{H}_{\mathbb{K}}}$. Here, $\mathcal{H}_{\mathbb{K}}$ is the Reproducing Kernel Hilbert Space (RKHS) associated with $\mathbb{K}$, and $\langle\cdot,\cdot\rangle_{\mathcal{H}{\mathbb{K}}}$ denotes the inner product in $\mathcal{H}_{\mathbb{K}}$. We assume that the feature mappings are bounded, i.e., $\|\phi(x)\|_{\mathcal{H}_{\mathbb{K}}}\leq R,\forall x\in\chi$ for some $R>0$.
Using this setup, we focus on an arbitrary fixed data point $x_{i}$ and compute its similarity with other data points $x_{j}(j\neq i)$ using the kernel function:
$S=\mathbb{K}\left(x_{i},x_{j}\right)=\left\langle\phi\left(x_{i}\right),\phi\left(x_{j}\right)\right\rangle_{\mathcal{H}_{\mathbb{K}}}$ (1)
Thus, we define the function $f$ as $f\left(x_{j}\right)=\left\langle w,\phi\left(x_{j}\right)\right\rangle_{\mathcal{H}_{\mathbb{K}}}$ where $w=\phi\left(x_{i}\right)$. Since $w=\phi\left(x_{i}\right)$ and $\left\|\phi\left(x_{i}\right)\right\|_{\mathcal{H}_{\mathbb{K}}}\leq R$, it follows that $\|w\|_{\mathcal{H}_{\mathbb{K}}}\leq R$. So the function $f$ representing the similarity between $x_{i}$ and any other point $x_{j}$ can be formulated as:
$f\left(x_{j}\right)=\left\langle\phi\left(x_{i}\right),\phi\left(x_{j}\right)\right\rangle_{\mathcal{H}_{\mathbb{K}}}=\mathbb{K}\left(x_{i},x_{j}\right)=S_{ij}.$ (2)
Since we are dealing with a ranking problem, we use the hinge loss function for pairs $\left(x_{j},x_{k}\right)(j,k\neq i)$ :
$\ell_{\text{hinge }}\left(f,x_{j},x_{k}\right)=\max\left(0,1-s_{jk}^{*}\left(f\left(x_{j}\right)-f\left(x_{k}\right)\right)\right),$ (3)
where $s_{jk}^{*}=\operatorname{sign}\left(f^{*}\left(x_{j}\right)-f^{*}\left(x_{k}\right)\right)$, and $f^{*}$ represents the ground-truth ranking function.
Thus, the expected loss over all pairs is given by:
$L(f)=\mathbb{E}_{\left(x_{j},x_{k}\right)}\left[\ell_{\text{hinge }}\left(f,x_{j},x_{k}\right)\right],$ (4)
and the empirical loss over the sample $S$ is given by:
$\hat{L}(f)=\frac{2}{(n-1)(n-2)}\sum_{\begin{subarray}{c}j<k\\ j,k\neq i\end{subarray}}\ell_{\text{hinge }}\left(f,x_{j},x_{k}\right),$ (5)
since there are $\frac{(n-1)(n-2)}{2}$ pairs among the $n-1$ data points excluding $x_{i}$. To assess how well the learned function $f$ generalizes to test data, we derive a generalization bound using Rademacher complexity. We define the function class for pairs by:
$\mathcal{F}_{\text{pair }}=\left\{f\left(x_{j},x_{k}\right)=f\left(x_{j}\right)-f\left(x_{k}\right)\right\}.$ (6)
which represents the similarity difference between $x_{j}$ and $x_{k}$ given the reference graph $x_{i}$.
Given that $f(x)=\langle w,\phi(x)\rangle$ with $\|w\|\leq R$ and $\|\phi(x)\|\leq R$, we can bound $f\left(x_{j},x_{k}\right)$ by:
$\left|f\left(x_{j},x_{k}\right)\right|=\left|f\left(x_{j}\right)-f\left(x_{k}\right)\right|\leq\left|f\left(x_{j}\right)\right|+\left|f\left(x_{k}\right)\right|\leq 2R^{2},$ (7)
since $|f(x)|=|\langle w,\phi(x)\rangle|\leq\|w\|\cdot\|\phi(x)\|\leq R\cdot R=R^{2}$.
Given $f\left(x_{j},x_{k}\right)$ takes values in the range $\left[-2R^{2},2R^{2}\right]$. Therefore, following the PAC-learning framework in [Mohri, 2018], we derive the following generalization bound. For any $\delta>0$, with probability at least $1-\delta$ over a sample $S$ of size $m$ drawn, the following holds for all $f\in$ $\mathcal{F}_{\text{pair }}:$
$L(f)\leq\hat{L}(f)+2\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)+M\sqrt{\frac{\ln(1/\delta)}{2m}},$ (8)
where:
  • ^$\hat{\mathfrak{R}}_{m}\left(\mathcal{F}_{\text{pair }}\right)$ is the empirical Rademacher complexity of $\mathcal{F}_{\text{pair }}$.
  • $M=1+2R^{2}$ is an upper bound on the hinge loss function $\ell_{\text{hinge }}$.
  • $m=\frac{(n-1)(n-2)}{2}$ is the number of pairs.
  • $\delta$ is the confidence level.
It is observed that as $m$ increases, the term $\sqrt{\frac{\ln(1/\delta)}{2m}}$ decreases at a rate proportional to $\sqrt{1/m}$. Next, we proceed by proving that the Rademacher complexity is bounded and decreases as $m$ increases.
Theorem A.1
Assuming $\|\phi(x)\|_{\mathcal{H}_{K}}\leq R$ for all $x\in\chi$, the empirical Rademacher complexity of the function class $\mathcal{F}_{\text{pair}}$ can be bounded as:
^$\hat{\mathfrak{R}}_{S}(\mathcal{F}_{\text{pair}})\leq\frac{2R^{2}}{\sqrt{m}}.$ (9)
Proof 1
Based on the definition of Rademacher complexity, for function class $\mathcal{F}_{\text{pair }}$ with a sample size $m$, the complexity ^$\hat{\mathfrak{R}}_{S}(\mathcal{F}_{\text{pair}})$ is defined as:
^$\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)=\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{F}_{\text{pair }}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}f\left(x_{j_{i}},x_{k_{i}}\right)\right],$ (10)
where $\sigma=\left(\sigma_{1},\ldots,\sigma_{m}\right)$ are independent Rademacher variables (i.e., each $\sigma_{k}$ takes the value +1 or -1 with equal probability), and $\left\{\left(x_{j_{i}},x_{k_{i}}\right)\right\}_{i=1}^{m}$ are pairs sampled from the data.
We can express ^$\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)$ as:
^$\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)=\frac{1}{m}\mathbb{E}_{\sigma}\left[\sup_{\|w\|\leq R}\left\langle w,\sum_{i=1}^{m}\sigma_{i}\left(\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right)\right\rangle\right].$ (11)
Applying the Cauchy-Schwarz inequality:
$\langle w,v\rangle_{\mathcal{H}_{\mathbb{K}}}\leq\|w\|_{\mathcal{H}_{\mathbb{K}}}\|v\|_{\mathcal{H}_{\mathbb{K}}}\leq R\|v\|_{\mathcal{H}_{\mathbb{K}}}$ (12)
we have:
^$\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)\leq\frac{1}{m}R\cdot\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{m}\sigma_{i}\left(\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right)\right\|\right].$ (13)
Using Jensen’s inequality in the form $\mathbb{E}[\|X\|]\leq\left(\mathbb{E}\left[\|X\|^{2}\right]\right)^{1/2}$, we obtain:
^$\hat{\mathfrak{R}}_{S}(\mathcal{F}_{\text{pair }})\leq\frac{R}{m}\left(\mathbb{E}_{\sigma}\left[\left\|\sum_{k=1}^{m}\sigma_{k}\left(\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right)\right\|_{\mathcal{H}_{K}}^{2}\right]\right)^{1/2}$ (14)
Since $\sigma_{i}$ are independent Rademacher variables, we have $\mathbb{E}\sigma\left[\sigma_{i}^{2}\right]=1$. Therefore:
$\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{m}\sigma_{i}\left(\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right)\right\|^{2}\right]$ $\leq\sum_{i=1}^{m}\mathbb{E}_{\sigma}\left[\sigma_{i}^{2}\right]\left\|\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right\|^{2}$ (15)
$=\sum_{i=1}^{m}\left\|\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right\|^{2}.$
Given that $\|\phi(y)\|_{\mathcal{H}_{\mathbb{K}}}\leq R$ , we have:
$\left\|\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right\|^{2}\leq(2R)^{2}=4R^{2}.$ (16)
Therefore,
$\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{m}\sigma_{i}\left(\phi\left(x_{j_{i}}\right)-\phi\left(x_{k_{i}}\right)\right)\right\|\right]\leq\sqrt{m\cdot 4R^{2}}=2R\sqrt{m}.$ (17)
Thus
^$\hat{\mathfrak{R}}_{S}\left(\mathcal{F}_{\text{pair }}\right)\leq\frac{1}{m}R\cdot 2R\sqrt{m}=\frac{2R^{2}}{m}\sqrt{m}=\frac{2R^{2}}{\sqrt{m}}.$ (18)
Therefore, the empirical Rademacher complexity decreases with $m$.